# A Lensing-Band Approach to Spacetime Constraints

### 11 Dec 2023 v1

@@ Abstract
General relativity’s prediction that all black holes share identical properties, irrespective of their size, can now be empirically tested using electromagnetic observations of supermassive black holes and gravitational waves from mergers of stellar-mass black holes. In this work, we focus on the electromagnetic side of this test and quantify the constraining power of very-long-baseline interferometry (VLBI). We demonstrate how to use lensing bands—annular regions on the observer’s screen surrounding the critical curve—to constrain the underlying spacetime geometry. Contingent upon a detection of a lensed VLBI feature, the resulting lensing-band framework allows us to exclude spacetimes for which said feature cannot arise from geodesics that traversed the equatorial plane more than once. Focusing on the first lensed image and tests of black-hole uniqueness, we employ a parametrized spacetime as a case study. We find that resolving geometric information that goes beyond the apparent size of the critical curve has the potential to lift degeneracies between different spacetime parameters. Our work thereby quantifies a conservative estimate of the constraining power of VLBI measurements and contributes to a larger effort to simultaneously constrain geometry and astrophysics.
@@ I Introduction
Do all astrophysical black holes source (up to rescaling) the same exterior gravitational field? In general relativity (GR), the answer is ‘yes’: uniqueness theorems imply that (i) the vacuum exterior of any spherically symmetric source is given by the Schwarzschild solution [1]; (ii) the vacuum exterior of a stationary axisymmetric (as well as uncharged and non-degenerate) black-hole horizon is given by the subextremal Kerr solution [2]; and (iii) the leading-order asymptotics of any stationary and axisymmetric source are unique and agree with the asymptotics of the Kerr solution [3]. Taken together, these three theorems guarantee that the vacuum exterior of stationary solutions in GR reproduces the Newtonian limit, allow us to uniquely identify the (asymptotic quantities) mass $M$ and angular momentum $J$, and fix the first parametrized post-Newtonian (PPN) parameters to $\beta_{\text{GR}}=1$ and $\gamma_{\text{GR}}=1$ [4, 5, 6]. Physically, the latter two PPN parameters respectively quantify the spatial curvature which is “generated” per unit rest mass and the degree of “nonlinearity” arising in the superposition law for gravity [7].
Figure 1: Different types of observations classified by their source mass $M_{\text{source}}$ (in units of Planck mass $M_{\text{Planck}}$) and their curvature scale $\xi$ (in units of inverse Planck area ${\ell_{\text{Planck}}}^{-2}$). The boundary of the gray region indicates the formation of a horizon (according to GR). The vertical lines indicate increasing radial distance to the source, along which light-rays and test masses (small dots denote the planets of the Solar System and Pluto) probe the background geometry.
@@
The above issue of disentangling geometry and astrophysics (at least in part) arises from the following dichotomy. On the one hand, the critical curve itself depends only on the geometry [24]. On the other hand, the full observable image will clearly depend on the astrophysical source. Subsequent lensed images approximate the critical curve exponentially well but carry some residual astrophysical dependence [41]. Recent progress towards quantifying this dependence is based on a precise theoretical characterization of gravitational lensing in the Kerr spacetime [27, 28, 42, 23]. We will refer to the resulting definition of an ($n$-fold) lensed image as the $n$-th photon ring.
@@
Previous works beyond GR have mostly relied on the assumption of black-hole uniqueness and have then explored deviations which are, nonetheless, unconstrained by other observations (see, for example, Refs. [46, 47, 48, 49]). This assumption effectively collapses the entire parameter space in Fig. 1 and assumes that gravity acts universally. There are several reasons to, nevertheless, remain agnostic. First, GR itself produces event horizons which can be reached by either going to sufficiently large curvature (stellar-mass black holes) or to sufficiently large source mass (AGNs). Similarly, deviations from GR could be tied to nonlinearities [50] or even to nonlocal quantities. Second, theories beyond GR can have multiple (stable) branches of black-hole solutions, see e.g., [51, 52]. Different mass scales may be prone to end up in one or another branch. Indeed, the formation history of AGNs is not well understood [53, 54]. Third, and maybe most importantly, whenever we are granted new observational opportunities, it is good practice to best not be guided by any theoretical prejudice.
Given the above status of the field, our work is motivated by three objectives. Our first objective is to provide a generic framework to constrain spacetime parameters, including the ones related to black-hole uniqueness, with VLBI observations. Such uniqueness tests for AGN have been recently been considered in spherical symmetry in Ref. [55]. Here, we provide the means to extend them beyond spherical symmetry. Our second objective is to quantify the constraining power of the first lensed image beyond null tests [45], and extend this work to deviations from the Kerr paradigm. Our third objective is to disentangle the above constraints on the underlying spacetime geometry from uncertainty about the astrophysical accretion process. Previous studies have attempted to tackle this issue by large computation-intensive joint-inference [56, 48, 57], by calibration factors [46, 58, 47, 33, 59, 60], or by means of the second lensing band [49]. Here, we extend the latter approach and propose a generic scheme which circumvents modelling or calibrating the impact of astrophysical uncertainties.
@@
The rest of this paper is organized as follows. In  Section II, we present the framework by reviewing the geometric concept of lensing bands (Section II.1) and discussing some observational avenues to extract a lensed emission region from VLBI measurements (Section II.2). With these two inputs, we present a generic scheme to translate a confident observation of lensed emission in the image plane into constraints on the underlying geometry. Since in the present work we are not using real data for the VLBI feature, we will refer to all quantitative statements as “projected constraints.” Our results aim to quantify the capability of VLBI observations to impose constraints. As a proof of concept, we then focus on VLBI observations of ${}^{\ast}$ [13, 39]. In Section III, we benchmark the framework by recovering the Kerr parameters (mass and spin) and the inclination to the source. In Section IV we apply the lensing-band framework to a parametric deviation of the Kerr metric, whose parameters are associated with black-hole non-uniqueness. Projected constraints on specific spacetimes can be obtained by expressing said spacetime within the parametrization and comparing the respective coefficients. We summarise our results and discuss the prospects of future work in Sec.V. Throughout the paper, we use geometric units in which $G_{\rm N}=c=1$, and the (,+,+,+)(-,+,+,+)( - , + , + , + ) metric signature.
@@ II The lensing-band framework
If one defines the “photon ring” as the lensed image associated with the above lensing-band definition, and assumes that a VLBI measurement has detected said photon ring, then the above statement can be rephrased as:
@@
In Section II.1, we review several available definitions of the term “lensed” image or “($n$-fold) lensed” image. We compare these definitions and, in particular, review the henceforth-used notion of equatorial lensing bands. In Section II.2, we briefly review some observational avenues toward confidently detecting the associated VLBI feature. Relying only on these two inputs, in Section II.3, we present an algorithm for excluding parameter regions in families of spacetimes such as the Kerr spacetime.
@@ II.1 On the definition of a lensed image
In general one may relate strong lensing to the deflection angle between the ingoing and the outgoing portion of the ray (see, for example, Ref. [63] for how to obtain the deflection angle in the Kerr spacetime).
@@ II.1.1 Black hole lensing bands
In particular, the $0^{\text{th}}$-order lensing band then corresponds to all light rays that reach the observer’s screen with a cut-out portion corresponding to the apparent location of the event horizon in the observer’s screen. Therefore, the $0^{\text{th}}$-order lensing band covers the entire image plane apart from the direct image of the event horizon, commonly referred to as the inner shadow [29, 30, 31], and corresponds to the direct, and therefore weakly lensed, image of the accretion structure. In contrast, already the $1^{\text{st}}$-order lensing band covers only a finite region within the image plane. Its inner (outer) boundary corresponds to rays originating from the location of the event horizon (from equatorial radial infinity) and arriving at the screen after crossing the equatorial plane once more. Each successive higher-order lensing band is contained within the former and, for $n\rightarrow\infty$, the $n^{\text{th}}$-order lensing band exponentially converges to the critical curve  [24].
@@ II.1.2 Alternative definitions
Several other (sometimes closely related) geometric definitions of lensing bands have been discussed in the literature, see, e.g., [65, Fig. 1] or [49, App. A]. In particular, this includes a definition via turning points [23] in the polar coordinate $\theta$: one may alternatively define the $m^{\text{th}}$-order lensing band as the image region in which incident geodesics exhibit $m+1$ turning points in $\theta$, i.e., for which the condition $d\cos(\theta)/d\lambda=0$ (with $\lambda$ denoting the affine parameter along the geodesic) is met $m+1$ times along the geodesic. We will refer to this as the “turning-point definition” of the lensing bands.
@@ II.1.3 Numerical approximation via bisection
Due to their simple geometric definition, the effort to obtain lensing bands is comparable to the effort to obtain the critical curve.
For any algebraically special spacetime in which geodesic motion is separable, we expect that the lensing bands can thus be expressed by closed-form one-dimensional integrals, see Ref. [44, App. A] for the Kerr case. In the non-separable case, each lensing-band boundary can be obtained by numerical bisection. In complete analogy to numerically obtaining the critical curve, this amounts to bisecting while numerically solving the geodesic equation, i.e., a coupled system of ODEs.
@@ II.2 On the observation of a lensed emission region
In the following Sections II.2.1, II.2.2 and II.2.3, we briefly review three current avenues to observationally extract a lensed emission region from VLBI data. In Section II.2.4, we then model a specific lensed emission region with which we obtain the projected constraints presented in the remainder of this paper.
@@ II.2.1 A hybrid imaging algorithm
One way to isolate evidence for a lensed image from foreground emission is hybrid imaging. Hybrid imaging combines rasterized image reconstruction with the specific modeling of expected features. For the data obtained in the 2017 EHT observation run of ${}^{\ast}$ [13], hybrid imaging algorithms [67, 68] indicated a Bayesian preference for a model fit upon inclusion of an expected thin-ring feature over a model fit without said thin-ring feature [39]. The authors interpret these results as first evidence of (with the definitions of the present work) $n=1$ lensed emission. We refer the reader to Refs. [40, 69] for other analyses where no evidence for the presence of lensed emission was found.
@@ II.2.2 An interferometric diameter inferred from the visibility amplitude
An image consisting of nested rings, as expected from GR, produces a cascade of damped oscillations on progressively longer baselines [23]. Such decomposition is displayed, for example, in Fig. 1 of Ref. [45] for $0\leq n\leq 2$. Consequently, the visibility amplitude of such an image will display a ringing pattern whose period of oscillation is related to a characteristic length/diameter in the image domain, modulated by an envelope whose specific fall-off is dictated by the ring thickness [23]. Due to these properties in the visibility domain, a space-based VLBI observation can provide evidence for a lensed emission region within the $\text{LB}^{n\geq 1}_{\text{VLBI}}$, if a ringing pattern in the visibility amplitude is detected. By sufficiently long baselines, we mean that, since the $n=0$ image is also ringlike, the baseline should be long enough to provide evidence of a change in the slope of the envelope of the interferometric signature [23]. The location, in baseline length, where the direct image ($n=0$) stops dominating the visibility domain depends on the width of the $n=0$ image itself, which is highly dependent on the accretion structure surrounding the black hole [41, 45].
@@ II.2.3 A phase shift on polarimetric images
Such a rotationally-symmetric polarization structure is expected to arise from the accretion flow of magnetically arrested disks (MAD) which are currently preferred by observations of ${}^{\ast}$[72]. At moderate (near-face-on) inclination, the above polarimetric signature has been confirmed in simulated interferometric observations [71]. In particular, the authors construct a gain-robust interferometric quantity and demonstrate that future observations of ${}^{\ast}$ have the potential to detect this polarization signature.
@@ II.2.4 Concrete implementation of an input VLBI feature
Irrespective of which of the above avenues the reader deems most promising, we demonstrate the potential of such an observation to rule out parameter ranges in an underlying family of spacetime geometries such as the Kerr spacetime. As a concrete proof of concept, we assume a lensed emission region $\text{LB}^{n\geq 1}_{\text{VLBI}}$ which matches the (upper prior limit of the) narrow-ring feature reported in Ref. [39], i.e., a geometrically circular ring of radius $\theta_{\text{ring}}=21.7\pm 0.1\,\mu\text{as}$ and fractional width $\psi=0.05$. We neglect the error in $\theta_{\text{ring}}$, since it is a subleading contribution. A graphic representation of the resulting lensed emission region $\text{LB}^{n\geq 1}_{\text{VLBI}}$ is given in Fig. 3.
We emphasize once more that the posteriors in [39] are insufficient to constrain the width or shape of the ring-like VLBI feature. In the following, we implicitly assume that joint data analysis of future observations, see, for instance Refs. [73, 74], can confidently associate such a thin-ring VLBI feature to lensed emission.
@@
$\arctan(\alpha)=\theta_{M}\equiv\frac{MG_{N}}{D\,c^{2}}\;.$ (1)
In all statistical inferences, we assume $D=16.8\pm 0.8\,\text{Mpc}$ as inferred in [75, App. I], cf. also [76, 77, 78] for references to the original measurements. For our exploratory analysis, we neglect the error on $D$ as subleading, recognizing that any error in $D$ fully degenerates with the inferred mass $M$.

### 05 Mar 2024 v2

@@ Abstract
General relativity’s prediction that all black holes are described by the Kerr metric, irrespective of their size, can now be empirically tested using electromagnetic observations of supermassive black holes and gravitational waves from mergers of stellar-mass black holes. In this work, we focus on the electromagnetic side of this test and quantify the constraining power of very-long-baseline-interferometry (VLBI) observations of emission generated by hot gas surrounding supermassive black holes. We demonstrate how to use lensing bands—annular regions on the observer’s screen surrounding the critical curve—to constrain the underlying spacetime geometry. Contingent upon a detection of a lensed VLBI feature, the resulting lensing-band framework allows us to exclude spacetimes for which said feature cannot arise from geodesics that traversed the equatorial plane more than once. Focusing on the first indirect image and tests of black-hole uniqueness, we employ a parametrized spacetime as a case study. We find that resolving geometric information that goes beyond the apparent size of the critical curve has the potential to lift degeneracies between different spacetime parameters. Our work thereby quantifies a conservative estimate of the constraining power of VLBI measurements and contributes to a larger effort to simultaneously constrain geometry and astrophysics.
@@ I Introduction
Do all astrophysical black holes source (up to rescaling) the same exterior gravitational field? In general relativity (GR), the answer is ‘yes’: uniqueness theorems imply that (i) the vacuum exterior of any spherically symmetric source is given by the Schwarzschild solution [1, 2]; (ii) the vacuum exterior of a stationary axisymmetric (as well as uncharged and non-degenerate) black-hole horizon is given by the subextremal Kerr solution [3]; and (iii) the leading-order asymptotics of any stationary and axisymmetric source are unique and agree with the asymptotics of the Kerr solution [4]. Taken together, these three theorems guarantee that the vacuum exterior of stationary solutions in GR reproduces the Newtonian limit, allow us to uniquely identify the (asymptotic quantities) mass $M$ and angular momentum $J$, and fix the first parametrized post-Newtonian (PPN) parameters to $\beta_{\text{GR}}=1$ and $\gamma_{\text{GR}}=1$ [5, 6, 7]. Physically, the latter two PPN parameters, respectively, quantify the spatial curvature which is “generated” per unit rest mass and the degree of “nonlinearity” arising in the superposition law for gravity [8].
Figure 1: Different types of observations classified by their source mass $M_{\text{source}}$ (in units of Planck mass $M_{\text{Planck}}$) and their curvature scale $\xi$ (in units of inverse Planck area ${\ell_{\text{Planck}}}^{-2}$). The boundary of the gray region indicates the formation of a horizon (according to GR). The vertical lines indicate increasing radial distance to the source, along which light rays and test masses (small dots denote the planets of the Solar System and Pluto) probe the background geometry.
@@
The above issue of disentangling geometry and astrophysics (at least in part) arises from the following dichotomy. On the one hand, the critical curve itself depends only on the geometry [25]. On the other hand, the full observable image will clearly depend on the astrophysical source. Subsequent lensed images approximate the critical curve exponentially well but carry some residual astrophysical dependence [42]. Recent progress towards quantifying this dependence is based on a precise theoretical characterization of gravitational lensing in the Kerr spacetime [28, 29, 43, 24]. We will refer to the resulting definition of an ($n$-fold) lensed image as the $n$-th photon ring, where the indexing $n+1$ is counting the maximum number of times an emitted photon that arrives to the observer plane has crossed the equatorial plane. Thus, for example, $n=0$ corresponds to the direct (lensed) image, i.e., photons that arrive to the observer’s screen and have crossed the equatorial plane at most once.
@@
Previous works beyond GR have mostly relied on the assumption of black-hole uniqueness and have then explored deviations which are, nonetheless, unconstrained by other observations (see, for example, Refs. [47, 48, 49, 50]). This assumption effectively collapses the entire parameter space in Fig. 1 and assumes that gravity acts universally. There are several reasons to, nevertheless, remain agnostic. First, GR itself produces event horizons which can be reached by either going to sufficiently large curvature (stellar-mass black holes) or to sufficiently large source mass (AGNs). Similarly, deviations from GR could be tied to nonlinearities [51] or even to nonlocal quantities. Second, theories beyond GR can have multiple (stable) branches of black-hole solutions, see e.g., Refs. [52, 53]. Different mass scales may be prone to end up in one or another branch. Indeed, the formation history of AGNs is not well understood [54, 55]. Third, and maybe most importantly, whenever we are granted new observational opportunities, it is good practice to best not be guided by any theoretical prejudice.
Given the above status of the field, our work is motivated by three objectives. Our first objective is to provide a generic framework to constrain spacetime parameters, including the ones related to black-hole uniqueness, with VLBI observations. Such uniqueness tests for AGN have been recently been considered in spherical symmetry in Ref. [56]. Here, we provide the means to extend them beyond spherical symmetry. Our second objective is to quantify the constraining power of the first indirect image beyond null tests [46], and extend this work to deviations from the Kerr paradigm. Our third objective is to disentangle the above constraints on the underlying spacetime geometry from uncertainty about the astrophysical accretion process. Previous studies have attempted to tackle this issue by large computation-intensive joint inference [57, 58, 59, 49, 60], by calibration factors [47, 61, 48, 34, 62, 63], or by means of the second lensing band [50]. Here, we extend the latter approach and propose a generic scheme which circumvents modelling or calibrating the impact of astrophysical uncertainties.
@@
The rest of this paper is organized as follows. In Sec. II, we present the framework by reviewing the geometric concept of lensing bands (Subsec. II.1) and discussing some observational avenues to extract a lensed emission region from VLBI measurements (Subsec. II.2). With these two inputs, we present a generic scheme to translate a confident observation of lensed emission in the image plane into constraints on the underlying geometry. Since in the present work we are not using real data for the VLBI feature, we will refer to all quantitative statements as “projected constraints.” Our results aim to quantify the capability of VLBI observations to impose constraints. As a proof of concept, we then focus on VLBI observations of ${}^{\ast}$ [14, 40]. In Sec. III, we benchmark the framework by recovering the Kerr parameters (mass and spin) and the inclination to the source. In Sec. IV we apply the lensing-band framework to a parametric deviation of the Kerr metric, whose parameters are associated with black-hole non-uniqueness. Projected constraints on specific spacetimes can be obtained by expressing said spacetime within the parametrization and comparing the respective coefficients. We summarize our results and discuss the prospects of future work in Sec. V. Throughout the paper, we use geometric units in which $G_{\rm N}=c=1$, and the (,+,+,+)(-,+,+,+)( - , + , + , + ) metric signature.
@@ II The lensing-band framework
If one defines the “photon ring” as the lensed image associated with the above lensing-band definition, and assumes that a VLBI measurement has detected said photon ring, then the above statement can be rephrased as follows:
@@
In Subsec. II.1, we review several available definitions of the term “lensed” image or “($n$-fold) lensed” image. We compare these definitions and, in particular, review the henceforth-used notion of equatorial lensing bands. In Subsec.II.2, we briefly review some observational avenues toward confidently detecting the associated VLBI feature. Relying only on these two inputs, in Subsec. II.3, we present an algorithm for excluding parameter regions in families of spacetimes such as the Kerr spacetime.
@@ II.1 On the definition of a lensed image
In general one may relate lensing to the deflection angle between the ingoing and the outgoing portion of the ray (see, for example, Ref. [66] for how to obtain the deflection angle in the Kerr spacetime).
@@ II.1.1 Black hole lensing bands
In particular, the zeroth-order lensing band then corresponds to all light rays that reach the observer’s screen with a cut-out portion corresponding to the apparent location of the event horizon in the observer’s screen. Therefore, the zeroth-order lensing band covers the entire image plane apart from the direct image of the event horizon, commonly referred to as the inner shadow [30, 31, 32], and corresponds to the direct, and therefore weakly lensed, image of the accretion structure. In contrast, already the first-order lensing band covers only a finite region within the image plane. Its inner (outer) boundary corresponds to rays originating from the location of the event horizon (from equatorial radial infinity) and arriving at the screen after crossing the equatorial plane once more. Each successive higher-order lensing band is contained within the former and, for $n\rightarrow\infty$, the $n^{\text{th}}$-order lensing band exponentially converges to the critical curve  [25].
@@ II.1.2 Alternative definitions
Several other (sometimes closely related) geometric definitions of lensing bands have been discussed in the literature, see, e.g., Refs. [23, Sec. 4.3], [68, Fig. 1], or [50, App. A]. In particular, this includes a definition via turning points [24] in the polar coordinate $\theta$: one may alternatively define the $m^{\text{th}}$-order lensing band as the image region in which incident geodesics exhibit $m+1$ turning points in $\theta$, i.e., for which the condition $d\cos(\theta)/d\lambda=0$ (with $\lambda$ denoting the affine parameter along the geodesic) is met $m+1$ times along the geodesic. We will refer to this as the “turning-point definition” of the lensing bands.
@@ II.1.3 Numerical approximation via bisection
Due to their simple geometric definition, the effort to obtain lensing bands is comparable to the effort to obtain the critical curve. For any algebraically special spacetime in which geodesic motion is separable, we expect that the lensing bands can thus be expressed by closed-form one-dimensional integrals, see Ref. [45, App. A] for the Kerr case. In the non-separable case, each lensing-band boundary can be obtained by numerical bisection. In complete analogy to numerically obtaining the critical curve, this amounts to bisecting while numerically solving the geodesic equation, i.e., a coupled system of ODEs.
@@ II.2 On the observation of a lensed emission region
In the following Subsecs. II.2.1 to II.2.3, we briefly review three current avenues to observationally extract a lensed emission region from VLBI data. In Subsec. II.2.4, we then model a specific lensed emission region with which we obtain the projected constraints presented in the remainder of this paper.
@@ II.2.1 A hybrid imaging algorithm
One way to isolate evidence for a lensed image from foreground emission is hybrid imaging. Hybrid imaging combines rasterized image reconstruction with the specific modeling of expected features. In Ref. [40], for the data obtained in the 2017 EHT observation run of ${}^{\ast}$ [14], hybrid imaging algorithms [70, 71] indicated a Bayesian preference for a model fit upon inclusion of an expected thin-ring feature over a model fit without said thin-ring feature. The authors of Ref. [40] interpreted these results as first evidence of (with the definitions of the present work) $n=1$ lensed emission. We refer the reader to Refs. [41, 72] for other analyses where no evidence for the presence of lensed emission was found.
@@ II.2.2 An interferometric diameter inferred from the visibility amplitude
An image consisting of nested rings, as expected from GR, produces a cascade of damped oscillations on progressively longer baselines [24]. Such decomposition is displayed, for example, in Fig. 1 of Ref. [46] for $0\leq n\leq 2$. Consequently, the visibility amplitude of such an image will display a ringing pattern whose period of oscillation is related to a characteristic length/diameter in the image domain, modulated by an envelope whose specific falloff is dictated by the ring thickness [24]. Due to these properties in the visibility domain, a space-based VLBI observation can provide evidence for a lensed emission region within the $\text{LB}^{n\geq 1}_{\text{VLBI}}$, if a ringing pattern in the visibility amplitude is detected. By sufficiently long baselines, we mean that, since the $n=0$ image is also ringlike, the baseline should be long enough to provide evidence of a change in the slope of the envelope of the interferometric signature [24]. The location, in baseline length, where the direct image ($n=0$) stops dominating the visibility domain depends on the width of the $n=0$ image itself, which is highly dependent on the accretion structure surrounding the black hole [42, 46].
@@ II.2.3 A phase shift on polarimetric images
Such a rotationally-symmetric polarization structure is expected to arise from the accretion flow of magnetically arrested disks which are currently preferred by observations of ${}^{\ast}$ [75]. At moderate (near-face-on) inclination, the above polarimetric signature has been confirmed in simulated interferometric observations [74]. In particular, the authors construct a gain-robust interferometric quantity and demonstrate that future observations of ${}^{\ast}$ have the potential to detect this polarization signature.
@@ II.2.4 Concrete implementation of an input VLBI feature
Irrespective of which of the above avenues the reader deems most promising, we demonstrate the potential of such an observation to rule out parameter ranges in an underlying family of spacetime geometries such as the Kerr spacetime. As a concrete proof of concept, we assume a lensed emission region $\text{LB}^{n\geq 1}_{\text{VLBI}}$ which matches the (upper prior limit of the) narrow-ring feature reported in Ref. [40], i.e., a geometrically circular ring of radius $\theta_{\text{ring}}=21.7\pm 0.1\,\mu\text{as}$ and fractional width $\psi=0.05$. We neglect the error in the radius $\theta_{\text{ring}}$, since it is a subleading contribution. A graphic representation of the resulting lensed emission region $\text{LB}^{n\geq 1}_{\text{VLBI}}$ is given in Fig. 3.
We emphasize once more that the posteriors in Ref. [40] are insufficient to constrain the width or shape of the ring-like VLBI feature. In the following, we implicitly assume that joint data analysis of future observations, see, for instance Refs. [76, 77], can confidently associate such a thin-ring VLBI feature to lensed emission.
@@
$\tan(\alpha)\equiv\frac{MG_{N}}{D\,c^{2}}\;.$ (1)
In all statistical inferences, we assume $D=16.8\pm 0.8\,\text{Mpc}$ as inferred in Ref. [78, App. I], cf. also Refs. [79, 80, 81] for the original measurements. For our exploratory analysis, we neglect the error on $D$ as subleading, recognizing that any error in $D$ fully degenerates with the inferred mass $M$.